calculate the perimeter of the rectangle, with sides of 30m and 15m
Please show me all the calculations/Solution from start to end

Found 3 solutions by josgarithmetic, Theo, ikleyn:
Answer by josgarithmetic(39617)

"sides"?

If you have a rectangle with dimensions x and y, then the perimeter is 2x+2y.

Answer by Theo(13342)

p = perimeter of the rectangle.
x = the length
y = the width.
p = 2 * (x + y)
that means the perimeters is equal to 2 times the sum of the length and the width.
x = 30m
y = 15m
p = 2 * 45m = 90m
here's a reference on perimeter of a rectangle.
https://www.varsitytutors.com/basic_geometry-help/how-to-find-the-perimeter-of-a-rectangle
the references uses l for the length and w for the width.
my formula would then be:
p = perimeter of the rectangle
l = length
w = width
p = 2 * (l + w)
this is the same as p = 2l + 2w.
i just factored out the 2.


Answer by ikleyn(52781)

.


Add the lengths of all 4 sides


    30 + 15 + 30 + 15 = 90 m.     ANSWER

Solved.

--------------

What can be easier than that ?